# Energy Matching: Unifying Flow Matching and Energy-Based Models for Generative Modeling

### v3

@@ Abstract
The most widely used generative models map noise and data distributions by matching flows or scores. However, they struggle to incorporate partial observations and additional priors—something energy-based models (EBMs) handle elegantly by simply adding corresponding scalar energy terms. We address this issue by proposing Energy Matching, a framework that endows flow-based approaches with the flexibility of EBMs. Far from the data manifold, samples move along curl-free, optimal transport paths from noise to data. As they approach the data manifold, an entropic energy term guides the system into a Boltzmann equilibrium distribution, explicitly capturing the underlying likelihood structure of the data. We parameterize this dynamic with a single time-independent scalar field, which serves as both a powerful generator and a flexible prior for effective regularization of inverse problems. Our method substantially outperforms existing EBMs on CIFAR-10 and ImageNet generation in terms of fidelity, while retaining simulation-free training of transport-based approaches away from the data manifold. Furthermore, we leverage the method’s flexibility to introduce an interaction energy that supports diverse mode exploration, which we demonstrate in a controlled protein-generation setting. Our approach focuses on learning a scalar potential energy—without time-conditioning, auxiliary generators, or additional networks—which marks a significant departure from recent EBM methods. We believe that this simplified framework significantly advances EBMs capabilities and paves the way for their wider adoption in generative modeling across diverse domains.
@@ 1 Introduction
Many generative models learn to map from a simple, easy-to-sample distribution, such as a Gaussian, to a desired data distribution. They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Tong et al., 2023) - or through iterative noising and denoising schemes, such as in diffusion models (Ho et al., 2020; Song et al., 2021). In addition to being highly effective in sample generation, diffusion- and flow-based models have also been used as priors to regularize poorly posed inverse problems (Chung et al., 2023; Mardani et al., 2024; Ben-Hamu et al., 2024). However, these models do not explicitly capture the unconditional data score and instead model the score of smoothed manifolds at different noise levels. The measurement likelihood, on the other hand, is not tractable on these noised manifolds. As a result, existing approaches repeatedly shuttle between noised and data distributions, leading to crude approximations of complex, intractable terms Daras et al. (2024). For example, DPS (Chung et al., 2023) approximates an intractable integral using a single sample. More recently, D-Flow (Ben-Hamu et al., 2024) optimizes initial noise by differentiating through the simulated trajectory. To the best of our knowledge, these models lack a direct way to navigate the data manifold in search of the optimal solution without repeatedly transitioning between noised and data distributions.
@@
Despite the theoretical elegance of using a single, time-independent scalar energy, practical EBMs have historically suffered from poor generation quality, falling short of the performance of diffusion or flow matching models. Traditional methods (Song and Kingma, 2021) for training EBMs, such as contrastive divergence via Markov chain Monte Carlo (MCMC) or local score-based approaches (Song and Ermon, 2019), often fail to adequately explore the energy landscape in high-dimensional spaces, leading to instabilities and mode collapse. Consequently, many methods resort to time-conditioned ensembles (Gao et al., 2021), hierarchical latent ensembles (Cui and Han, 2024), or combine EBMs with separate generator networks trained in cooperation (Guo et al., 2023; Zhang et al., 2024; Yoon et al., 2024), thereby requiring significantly higher parameter counts and training complexity. {mdframed}[hidealllines=true,backgroundcolor=blue!5]
@@
In this work, we propose Energy Matching, a two-regime training strategy that combines the strengths of EBMs and flow matching; see Figure 1.
When samples lie far from the data manifold, they are efficiently transported toward the data. Once near the data manifold, the flow transitions into Langevin steps governed by an internal energy component, enabling precise exploration of the Boltzmann-like density well around the data distribution. This straightforward approach produces a time-independent scalar energy field whose gradient both accelerates sampling and shapes the final density well—via a contrastive objective that directly learns the score at the data manifold—yet remains efficient and stable to train. Empirically, our method significantly outperforms existing EBMs on both CIFAR-10 and ImageNet generation in terms of fidelity, and compares favorably to flow-matching and diffusion models—without auxiliary generators or time-dependent EBM ensembles.
Our proposed process complements the advantages of flow matching with an explicit likelihood modeling, enabling traversal of the data manifold without repeatedly shuffling between noise and data distributions. This simplifies both inverse problem solving and controlled generations under a prior. In addition, to encourage diverse exploration of the data distribution, we showcase how repulsive interaction energies can be easily and effectively incorporated, with an application to conditional protein generation. Finally, we also showcase how analyzing the learned energy reveals insight on the LID of the data with fewer approximations than diffusion models. [^1] The code will be publicly released upon acceptance.
@@ 2 Energy matching
In this section, we show how a scalar potential $V(x)$ can simultaneously provide an optimal-transport-like flow from noise to data while also yielding a Boltzmann distribution that explicitly captures the unnormalized likelihood of the data.
@@
$W_{2}^{2}(\rho,\rho_{t})=\min_{\gamma\in\Gamma(\rho,\rho_{t})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\norm{x-x_{t}}^{2}\mathrm{d}\gamma(x,x_{t}),$ (2)
@@
Differently from most literature, we consider $\varepsilon(t)$ to be dependent on time and study the behavior of Equation 1 as $t\to\infty$. To fix the ideas, consider, for instance, a linear scheduling:
$\varepsilon(t)=\begin{cases}0,&0\leq t<\operatorname{\tau^{\ast}},\\ \varepsilon_{\max}\frac{t-\operatorname{\tau^{\ast}}}{1-\operatorname{\tau^{\ast}}},&\operatorname{\tau^{\ast}}\leq t\leq 1,\\ \varepsilon_{\max},&t\geq 1.\end{cases}$ (3)
@@
We follow the approach in Terpin et al. (2024) and study (1) at each time $t$ via its first-order optimality conditions (Lanzetti et al., 2024, 2025):
$\frac{1}{\Delta t}(x-y)+\nabla_{x}V_{\theta}(x)+\varepsilon(t)\nabla_{x}\log\rho_{t}(x)=0,\qquad(x,y)\in\mathrm{supp}(\gamma_{t})$ (4)
where $\gamma_{t}$ is an OT plan between the distributions $\rho_{t}$ and $\rho_{t+\Delta t}$ and $\mathrm{supp}(\gamma_{t})$ is the support of $\gamma_{t}$. That is, this condition has to hold for all pairs of points in the support of $\rho_{t}$ and $\rho_{t+\Delta t}$ that are coupled by OT. Intuitively, analyzing (4) provides us with two key insights:
@@
    For times $t<\tau^{*}$, $\varepsilon(t)=0$ and (4) becomes
@@
    Near the data manifold, which we aim at modeling with the equilibrium distribution $\rho_{\text{eq}}$ of (1), $\rho_{t}\approx\rho_{\text{eq}}$ and, thus, for $t\gg 1$, $x\approx y$ for all $(x,y)\in\mathrm{supp}(\gamma_{t})$. Then, we can simplify (4) as
    $\varepsilon_{\max}\nabla_{x}\log\rho_{\mathrm{eq}}(x)=\nabla_{x}V_{\theta}(x)\quad\Longrightarrow\quad\rho_{\mathrm{eq}}(x)\propto\exp\left(-\tfrac{V_{\theta}(x)}{\varepsilon_{\max}}\right).$
@@
By combining the long-range transport capability of flows with the local density modeling flexibility of EBMs, we achieve tractable sampling and explicit likelihood of samples $V_{\theta}(x)$; see Figure 1.
@@ 2.1 Training objectives
In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$, see Section 2.1.1), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase. Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach equilibrium (see Section 2.1.2).
@@ 2.1.1 Flow-like objective $\operatorname{\mathcal{L}_{\mathrm{OT}}}$
We begin by constructing a global velocity field $\nabla_{x}V_{\theta}(x)$ that carries noise samples $\{x_{0}\}$ to data samples $\{x_{\mathrm{data}}\}$ with minimal detours. For this, we consider geodesics in the Wasserstein space (Ambrosio et al., 2008). Practically, we compute the OT coupling $\gamma^{*}$ between two uniform empirical probability distributions, one supported on a mini-batch $\{x_{\text{data}}^{(j)}\}_{j=1}^{B}$ of the data, and one supported on a set of noise samples with the same cardinality $B$. These samples are drawn from an easy-to-sample distribution; in our case, a Gaussian. Since the probability distributions are uniform and empirical with the same number of samples, a transport map $T$ is guaranteed to exist (Ambrosio et al., 2008).
@@
Depending on the approach used to compute the OT coupling, one may find an OT map or not. Similarly, if the set of noise samples has cardinality smaller or larger than $B$, the OT coupling will not be expressed by a map. In this case, one can adapt our algorithm by defining a threshold $\pi_{\mathrm{th}}$ and consider all the pairs $\left(x_{0}^{(j)},x_{\mathrm{data}}^{(i)}\right)$ with $\gamma^{\ast}(x_{0}^{(j)},x_{\mathrm{data}}^{(i)})>\pi_{\mathrm{th}}$. In our experiments, we used the POT solver (Flamary et al., 2021) and we did not notice benefits in a sample size other than $B$, in line with the observations in (Tong et al., 2023).
Then, for each data point $x_{\mathrm{data}}^{(i)}$ we define the interpolation $x_{t}^{(i)}=(1-t)T(x_{\mathrm{data}}^{(i)})+tx_{\mathrm{data}}^{(i)}$, which is a point along the geodesic (Ambrosio et al., 2008). The velocity of each $x_{t}^{(i)}$ is $x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$ (i.e., the samples move from the noise to the data distribution at constant speed) and, in this regime, we would like to have $-\nabla_{x}V_{\theta}(x_{t}^{(i)})\approx x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$. For this, we define the loss:
$\operatorname{\mathcal{L}_{\mathrm{OT}}}=\mathbb{E}^{\begin{subarray}{c}x_{\mathrm{data}}\in\mathcal{D}\\ t\sim U(0,1)\end{subarray}}\left[\norm{\nabla_{x}V_{\theta}(x_{t}^{(i)})+x_{\mathrm{data}}^{(i)}-T({x^{(i)}_{\mathrm{data}}})}^{2}\right].$
This objective can be interpreted as a flow-matching formulation under the assumption that the velocity field is both time-independent and given by the gradient of a scalar potential, thereby imposing a curl-free condition. This aligns naturally with OT, which also yields a velocity field of zero curl—any rotational component would add unnecessary distance to the flow and thus inflate the transport cost without benefit. Our experimental evidence adds to the recent study by (Sun et al., 2025), in which the authors observed that time-independent velocity fields can, under certain conditions, outperform time-dependent noise-conditioned fields in sample generation.
@@
2:for iteration $0,1,\ldots$ do
3:    Sample $\{x_{\mathrm{data}}^{(j)}\}_{j=1}^{B}\sim\mathcal{D}$triangleright Mini-batch of data samples
4:    Sample $\{x_{0}^{(i)}\}_{i=1}^{B}\sim\mathcal{N}(0,I)$triangleright Random Gaussian samples
5:    $T\leftarrow\mathrm{OTsolver}(\{x_{0}^{(i)}\},\{x_{\mathrm{data}}^{(j)}\})$triangleright Compute OT map
6:    Sample $t\sim U(0,1)$
7:    $\operatorname{\mathcal{L}_{\mathrm{OT}}}(\theta)\leftarrow\sum_{i}\|\nabla_{x}V_{\theta}(x_{t}^{(i)})+x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})\|^{2}$triangleright Loss function
8:    $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\text{OT}}(\theta)$triangleright Gradient update with learning rate $\alpha$
9:end for
10:return $\theta$$\triangleright$ Trained $\theta$

### v5

@@ Abstract
Current state-of-the-art generative models map noise to data distributions by matching flows or scores. A key limitation of these models is their inability to readily integrate available partial observations and additional priors. In contrast, energy-based models (EBMs) address this by incorporating corresponding scalar energy terms. Here, we propose Energy Matching, a framework that endows flow-based approaches with the flexibility of EBMs. Far from the data manifold, samples move from noise to data along irrotational, optimal transport paths. As they approach the data manifold, an entropic energy term guides the system into a Boltzmann equilibrium distribution, explicitly capturing the underlying likelihood structure of the data. We parameterize these dynamics with a single time-independent scalar field, which serves as both a powerful generator and a flexible prior for effective regularization of inverse problems. The present method substantially outperforms existing EBMs on CIFAR-10 and ImageNet generation in terms of fidelity, while retaining simulation-free training of transport-based approaches away from the data manifold. Furthermore, we leverage the flexibility of the method to introduce an interaction energy that supports the exploration of diverse modes, which we demonstrate in a controlled protein generation setting. This approach learns a scalar potential energy, without time conditioning, auxiliary generators, or additional networks, marking a significant departure from recent EBM methods. We believe this simplified yet rigorous formulation significantly advances EBMs capabilities and paves the way for their wider adoption in generative modeling in diverse domains.
@@ 1 Introduction
Generative models learn to map from a simple, easy-to-sample distribution, such as a Gaussian, to a desired data distribution. They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Albergo and Vanden-Eijnden, 2023)—or through iterative noising and denoising schemes, such as in diffusion models (Ho et al., 2020; Song et al., 2021). In addition to being highly effective in sample generation, diffusion- and flow-based models have also been used as priors to regularize poorly posed inverse problems (Chung et al., 2023; Mardani et al., 2024; Ben-Hamu et al., 2024). However, these models do not explicitly capture the unconditional data score and instead model the score of smoothed manifolds at different noise levels. The measurement likelihood, on the other hand, is not tractable on these noised manifolds. As a result, existing approaches repeatedly shuttle between noised and data distributions, leading to crude approximations of complex, intractable terms Daras et al. (2024). For example, DPS (Chung et al., 2023) approximates an intractable integral using a single sample. More recently, D-Flow (Ben-Hamu et al., 2024) optimizes initial noise by differentiating through the simulated trajectory. To the best of our knowledge, these models lack a direct way to navigate the data manifold in search of the optimal solution without repeatedly transitioning between noised and data distributions.
@@
Despite the theoretical elegance of using a single, time-independent scalar energy, practical EBMs have historically suffered from poor generation quality, falling short of the performance of diffusion or flow matching models. Traditional methods (Song and Kingma, 2021) for training EBMs, such as contrastive divergence via Markov chain Monte Carlo (MCMC) or local score-based approaches (Song and Ermon, 2019), often fail to adequately explore the energy landscape in high-dimensional spaces, leading to instabilities and mode collapse. Consequently, many methods resort to time-conditioned ensembles (Gao et al., 2021), hierarchical latent ensembles (Cui and Han, 2024), or combine EBMs with separate generator networks trained in cooperation (Guo et al., 2023; Zhang et al., 2024; Yoon et al., 2024), thereby requiring significantly higher parameter counts and training complexity.
@@
In this work, we propose Energy Matching, a two-regime training strategy that combines the strengths of EBMs and flow matching; see Figure˜1. When samples lie far from the data manifold, they are efficiently transported toward the data. Once near the data manifold, the flow transitions into Langevin steps governed by an internal energy component, enabling precise exploration of the Boltzmann-like density well around the data distribution. This straightforward approach produces a time-independent scalar energy field whose gradient both accelerates sampling and shapes the final density well—via a contrastive objective that directly learns the score at the data manifold—yet remains efficient and stable to train. Empirically, our method significantly outperforms existing EBMs on both CIFAR-10 and ImageNet generation in terms of fidelity, and compares favorably to flow-matching and diffusion models—without auxiliary generators or time-dependent EBM ensembles. Our proposed process complements the advantages of flow matching with an explicit likelihood modeling, enabling traversal of the data manifold without repeatedly shuffling between noise and data distributions. This simplifies both inverse problem solving and controlled generations under a prior. In addition, to encourage diverse exploration of the data distribution, we showcase how repulsive interaction energies can be easily and effectively incorporated, with an application to conditional protein generation. Finally, we also showcase how analyzing the learned energy reveals insight on the LID of the data with fewer approximations than diffusion models. [^1] Code repository: https://github.com/m1balcerak/EnergyMatching
@@ 2 Energy matching
In this section, we show how a scalar potential $V(x)$ can simultaneously provide an optimal-transport-like flow from noise to data while also yielding a Boltzmann distribution that explicitly captures the unnormalized log‑likelihood of the data.
@@
$W_{2}^{2}(\rho,\rho_{t})=\min_{\gamma\in\Gamma(\rho,\rho_{t})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\norm{x-y}^{2}\mathrm{d}\gamma(x,y),$ (2)
@@
Differently from most literature, we consider $\varepsilon(t)$ to be dependent on time and study the behavior of Equation˜1 as $t\to\infty$. To fix the ideas, consider, for instance, a linear scheduling:
$\varepsilon(t)=\begin{cases}0,&0\leq t<\operatorname{\tau^{\ast}},\\ \varepsilon_{\max}\frac{t-\operatorname{\tau^{\ast}}}{1-\operatorname{\tau^{\ast}}},&\operatorname{\tau^{\ast}}\leq t<1,\\ \varepsilon_{\max},&t\geq 1.\end{cases}$ (3)
@@
Following Terpin et al. (2024), we analyze (1) at each time $t$ via its first-order optimality conditions (Lanzetti et al., 2024, 2025). These conditions characterize the properties of the desired solution and thus represent the optimization goal:
$\frac{1}{\Delta t}(x-y)+\nabla_{x}V_{\theta}(x)+\varepsilon(t)\nabla_{x}\log\rho_{t+\Delta t}(x)=0,\qquad(x,y)\in\mathrm{supp}(\gamma_{t})$ (4)
where $\gamma_{t}$ is an OT plan between the distributions $\rho_{t+\Delta t}$ and $\rho_{t}$ and $\mathrm{supp}(\gamma_{t})$ is the support of $\gamma_{t}$. That is, this condition has to hold for all pairs of points in the support of $\rho_{t+\Delta t}$ and $\rho_{t}$ that are coupled by OT. Intuitively, analyzing (4) provides us with two key insights:
@@
    For times $t<\operatorname{\tau^{\ast}}$, $\varepsilon(t)=0$ and (4) becomes
@@
    Near the data manifold, which we aim at modeling with the equilibrium distribution $\rho_{\text{eq}}$ of (1), $\rho_{t+\Delta t}\approx\rho_{\text{eq}}$ and, thus, for $t\gg 1$, $x\approx y$ for all $(x,y)\in\mathrm{supp}(\gamma_{t})$. Then, we can simplify (4) as
    $\varepsilon_{\max}\nabla_{x}\log\rho_{\mathrm{eq}}(x)=-\nabla_{x}V_{\theta}(x)\quad\Longrightarrow\quad\rho_{\mathrm{eq}}(x)\propto\exp\left(-\tfrac{V_{\theta}(x)}{\varepsilon_{\max}}\right).$
@@
By combining the long-range transport capability of flows with the local density modeling flexibility of EBMs, we achieve tractable sampling and explicitly encode the unnormalized log-likelihood $V_{\theta}(x)/\varepsilon_{\max}$ of the underlying data distribution; see Figure˜1.
@@ 2.1 Training objectives
In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$, see Section˜2.1.1), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase. Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach the data manifold (i.e., the equilibrium distribution); see Section˜2.1.2.
@@ 2.1.1 Flow-like objective $\operatorname{\mathcal{L}_{\mathrm{OT}}}$
We begin by constructing a global velocity field $\nabla_{x}V_{\theta}(x)$ that carries noise samples $\{x_{0}\}$ to data samples $\{x_{\mathrm{data}}\}$ with minimal detours. For this, we consider geodesics in the Wasserstein space (Ambrosio et al., 2008). Practically, we compute the OT coupling $\gamma^{*}$ between two uniform empirical probability distributions, one supported on a mini-batch of the data, and one supported on a set of noise samples with the same cardinality. These samples are drawn from an easy-to-sample distribution; in our case, a Gaussian. Since the probability distributions are uniform and empirical with the same number of samples, a transport map $T$ is guaranteed to exist (Ambrosio et al., 2008).
@@
Depending on the method used to compute the OT coupling, an explicit OT map may or may not be obtained. Similarly, if the number of noise samples differs from the mini-batch size $B$, the resulting OT coupling generally will not correspond to a map. In this case, one can adapt the algorithm by defining a threshold $\pi_{\mathrm{th}}$ and considering all pairs $(x_{\mathrm{data}},x_{0})$ for which the coupling value satisfies $\gamma^{\ast}(x_{\mathrm{data}},x_{0})>\pi_{\mathrm{th}}$. In our experiments, we used the POT solver (Flamary et al., 2021) and did not observe benefits from using a sample size different from $B$, consistent with previous approaches (Tong et al., 2023).
Then, for each data point $x_{\mathrm{data}}$ we define the interpolation $x_{t}=(1-t)T(x_{\mathrm{data}})+tx_{\mathrm{data}}$, which is a point along the geodesic. The velocity of each $x_{t}$ is $x_{\mathrm{data}}-T(x_{\mathrm{data}})$ (i.e., the samples move from the noise to the data distribution at constant speed) and, in this regime, we would like to have $\nabla_{x}V_{\theta}(x_{t})\approx x_{\mathrm{data}}-T(x_{\mathrm{data}})$. For this, we define the loss:
$\operatorname{\mathcal{L}_{\mathrm{OT}}}=\mathbb{E}{\begin{subarray}{c}x_{\mathrm{data}}\in\mathcal{D}\\ t\sim U(0,\operatorname{\tau^{\ast}})\end{subarray}}\left[\norm{\nabla_{x}V_{\theta}(x_{t})+x_{\mathrm{data}}-T({x_{\mathrm{data}}})}^{2}\right].$
This objective can be interpreted as a flow-matching formulation under the assumption that the velocity field is both time-independent and given by the gradient of a scalar potential, thereby imposing an irrotational condition. This aligns naturally with OT, which also yields an irrotational velocity field—any rotational component would add unnecessary distance to the flow and thus inflate the transport cost without benefit. Our experimental evidence adds to the recent study by (Sun et al., 2025), in which the authors observed that time-independent velocity fields can, under certain conditions, outperform time-dependent noise-conditioned fields in sample generation.
@@
2:for iteration $n=0,1,\ldots$ do
3:  Sample mini-batch $\{x_{\mathrm{data},b}\}_{b=1}^{B}\sim\mathcal{D}$ $\triangleright$ Data samples
4:  Sample mini-batch $\{x_{0,b}\}_{b=1}^{B}\sim\mathcal{N}(0,I)$ $\triangleright$ Random Gaussian samples
5:  $T\leftarrow\mathrm{OTsolver}(\{x_{\mathrm{data},b}\},\{x_{0,b}\})$ $\triangleright$ Compute OT map
6:  Sample $\{t_{b}\}_{b=1}^{B}\sim U(0,\operatorname{\tau^{\ast}})$ $\triangleright$ Typically $\operatorname{\tau^{\ast}}=1$ for the warm-up
7:  Set interpolations $x_{t_{b}}\leftarrow(1-t_{b})\,T(x_{\mathrm{data},b})+t_{b}\,x_{\mathrm{data},b}$ $\triangleright$ Interpolation along geodesics
8:  $\operatorname{\mathcal{L}_{\mathrm{OT}}}(\theta)\leftarrow\sum_{b=1}^{B}\|\nabla_{x}V_{\theta}(x_{t_{b}})+x_{\mathrm{data},b}-T(x_{\mathrm{data},b})\|^{2}$ $\triangleright$ Loss function
9:  $\theta\leftarrow\theta-\alpha\nabla_{\theta}\operatorname{\mathcal{L}_{\mathrm{OT}}}(\theta)$ $\triangleright$ Gradient update with learning rate $\alpha$
10:end for
11:return $\theta\triangleright$ Trained $\theta$